OpenStudy (anonymous):

If f(x) = x − 4 and g(x) 1/2= x − 2, find f[g(−6)].

OpenStudy (bahrom7893):

Just do what we've been doing so far.

OpenStudy (anonymous):

i know it had a fraction and it threw me off big time im horrible in math

OpenStudy (bahrom7893):

g(x)/2 = x-2 2*g(x)/2 = 2*(x-2) g(x) = 2x-4

OpenStudy (bahrom7893):

Does that help?

OpenStudy (anonymous):

kinda hold on let me try and solve it

OpenStudy (anonymous):

is it 3 ?

OpenStudy (bahrom7893):

can you post the steps? Saves me time and effort to just look at your solution instead of solving it over, and also lets me know if and where you went wrong.

OpenStudy (anonymous):

g(f(x))=2((1/2)x-3)+5=x-6+5=x-1 g(f(4))=4-1=3 can u understand that..

OpenStudy (anonymous):

for the first part i solved it like this... f(x)=(1/2)x-3 f(4)=(1/2)4-3 f(4)=2-3 f(4)=-1 then g(x)=2x+5 g(f(4))=g(-1)=2(-1)+5=-2+5=3 g(f(4))=3